OpenStudy (anonymous):

you toss 3 coins.find the probability atleast 1 tail is tossed A. 7/8 B. 1/8 C. 3/4 D. 1/3

OpenStudy (anonymous):

@shannon342

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@shannon342 thats wrong

OpenStudy (anonymous):

it says d is incorrect

OpenStudy (anonymous):

@shannon342 also, don't give answers without at least explaining

OpenStudy (anonymous):

he probability that at least one tail will be tossed is 1.00 - 0.125 = 0.875 (which is the same probability that at least one head will be tossed.)

OpenStudy (anonymous):

ur right but that would = 7/8 cuz 7/8 = 0.875

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes that was my mistake

OpenStudy (anonymous):

yes, what you're suppose to do is 1/2 + 1/4 + 1/8 = 7/8

OpenStudy (anonymous):

oh that is ok:P

OpenStudy (anonymous):

thank u guys so much now i get to take my end of year exam!